OpenStudy (anonymous):

Find the midpoint between two points on a number line if one of the points is at -7, and the other point is at 12. 5 9.5 2.5 -2

OpenStudy (calculusxy):

M = midpoint x_1 and x_2 = the two points (numbers) on the number line \[M = \frac{ x_1 + x_2 }{ 2 }\]

OpenStudy (calculusxy):

Do you understand the formula?

OpenStudy (calculusxy):

@ltj12345.

OpenStudy (anonymous):

yes

OpenStudy (calculusxy):

So plug in the variables and then solve it. Tell me the answer once you get it so that I can confirm it with you.

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

5

OpenStudy (calculusxy):

Let's check : \[M = \frac{ x_1 + x_2 }{ 2 } = \frac{ -7 + 12 }{ 2 } = \frac{ 5 }{ 2 } = 2.5\]

OpenStudy (calculusxy):

There might've been something wrong when you figured out 5. Did you catch your mistake?

OpenStudy (anonymous):

yes

OpenStudy (calculusxy):

So your answer is 2.5 .

OpenStudy (anonymous):

thanks

OpenStudy (calculusxy):

Great job!

OpenStudy (calculusxy):

You're welcome :)